OpenStudy (anonymous):

express the tan(22.5) using the cos(45)

OpenStudy (freckles):

\[22.5=\frac{45}{2} \\ \tan(\frac{45}{2}) =?..\text{ use half angle identity }\]

OpenStudy (freckles):

you may also consider expressing sin(45) as cos(45) you can do this since sin(45)=cos(45) I'm only saying this since they want you to express your answer in terms of cos(45)

OpenStudy (anonymous):

I got -1-sqrt of 2 all over 1 so it seems right lol